APACHE POINT OBSERVATORY SDSS 2.5M OBSERVING LOG Monday January 1, 2007 (MJD 54102) ---=== OBSERVING TEAM ===--- Howard Brewington (Observer) Jian Ge (UF) P.C. Guo (UF) Craig Loomis (support) ---=== OBSERVING PLAN ===--- ET Science! ---=== OBSERVING SUMMARY ===--- Cloud, high humidity, and snow; we never opened. ---=== OBSERVING LOG ===--- Afternoon: ----------- During the telescope checkout, the telrun session stopped working after switching back from the mcp beyond the top and bottom balance check. I called Craig, but he could not find an obvious problem. So, Craig filed a a PR. Watcher is still unhappy about sp2's mechanicals, viz., "Not Connected." And, tccMon is still slow to start after issuing the tccNode command. Night: ----------- 00:07Z Clouds and high humidity at sunset, so we waited. 04:39Z Watcher reported that CCD-55 was upset about its vs+ voltage, which was 7.05 volts. This happened a few days ago during my dayShift checkout. 12:25Z Clouds, high humidity, and snow, I went to bed. ---=== IMAGING RUN SUMMARY ===--- Run Time Stripe Lambda Last Flavor Comments Start End Begin End Frame ------------------------------------------------------------------------- ---=== IMAGING RUN DETAILS ===--- ---=== SKIPPY RESULTS ===--- Run Frame nFrames stars muErr muRms nuErr nuRms rot az el --------------------------------------------------------------------------- ---=== LTMATCH RESULTS ===--- Run Field nFields alt az nGood rowMean rowSig colMean colSig rot ------------------------------------------------------------------------ ---=== SPECTROSCOPY DATA SUMMARY ===--- Summary Checked (y/n): QA Procedures Done (y/n): UT Exp Time flavor comment (S/N)^2 totals ========================================== b1 r1 b2 r2 ----- sequence 42285, plate -9999 ------- 21:41 42285 0.0 bias ---=== TELESCOPE OFFSETS AND SCALE I ===--- Time Instrument Az Alt Rot Scale pos offset pos offset pos offset ------------------------------------------------------------------------------ ---=== TELESCOPE OFFSETS AND SCALE II ===--- ---=== FOCUS LOG ===--- setmir piston Temp Wind Time Inst scale M1 M2 Foc Az Alt (C) MPH Dir filt fwhm ------------------------------------------------------------------------------ ---=== WEATHER LOG ===--- Wind Time Temp F Dewp F MPH Direction Dust DIMM Sky 20:50Z 35 17 4 240 (WSW) 142 - 21:22Z 35 18 17 170 (S) 142 - 21:52Z 34 15 11 147 (SSE) 99 - 22:22Z 33 17 17 152 (SSE) 113 - 22:54Z 31 18 23 155 (SSE) 155 - 23:26Z 30 17 17 155 (SSE) 164 - 23:58Z 29 18 17 127 (SE) 168 - 00:30Z 28 19 13 113 (ESE) 191 - 01:03Z 29 19 19 125 (SE) 258 - 01:59Z 27 20 16 124 (SE) 307 - 02:31Z 27 21 22 127 (SE) 243 - 03:04Z 27 21 14 114 (ESE) 284 - 03:36Z 27 21 16 112 (ESE) 257 - 04:08Z 27 21 18 124 (SE) 264 - 04:40Z 27 21 19 124 (SE) 249 - 09:37Z 28 23 15 132 (SE) 411 - 10:09Z 27 23 8 130 (SE) 437 - 10:41Z 27 23 3 169 (S) 447 - 11:14Z 26 23 3 121 (ESE) 407 - 11:46Z 27 23 1 182 (S) 419 - 12:18Z 27 23 0 22 (NNE) 404 - ---=== TELESCOPE STATUS ===--- 01:15Z Doors and louvers opened, fans on. 04:00Z Doors and louvers closed, fanns off. No interlocks bypassed. Status at: 12:20Z Telescope stowed at: 30,121,00 mount Instrument mounted: Cartridge 03 Counterweights at: 250 lbs LN2 autofill systems: Connected and turned on. 180L LN2 dewar scales: Spectro 108lbs, 22 psi Imager 211lbs, 22psi ---=== SOFTWARE USED ===--- IOP/SOP: v4_29_0 Watcher: v2_35_0 MCP: v6_3_2 TPM: tpm_v3_5_0 AstroDa: v15_13 TCC: TCC 2.7.2.1 August 6 2004 sdssProcedures: v2_21 SoS: v5_1_4 hoggPT: v1_6_9 plate-mapper: v4_3_1 ---=== MIRROR NUMBERS ===--- PRIMARY: -------- Scale: 1.000000 MIGS TONIGHT NOMINAL Axial A 24.6770 24.6730 Axial B 26.2720 26.2680 Axial C 27.9350 27.9320 Trans D 13.5820 13.5820 Lateral E 9.9695 9.9695 Lateral F 11.5570 11.5443 GALILS Commanded: 4600. -2000. 2650. -9300. -9800. -9800. Actual: 4600. -2000. 2650. -9300. -9800. -9800. SETMIR VALUES PriDesOrient: 0.00 0.00 0.00 0.00 0.00 PriOrient: 0.00 0.00 0.00 0.00 0.00 SECONDARY: ---------- Focus: 0.00 Air Temp.: 1.3 Alt.: 29.999596 MIGS TONIGHT NOMINAL Axial A 1.5510 1.5540 Axial B 1.1410 1.1390 Axial C 1.1930 1.1920 Trans D -0.0420 -0.0510 GALILS Commanded: 1614061. 1594704. 1543095. -10550. -5800. Actual: 1614050. 1594700. 1543100. -10550. -5800. SETMIR VALUES SecDesOrient: 1257.00 20.00 -10.00 -100.00 176.64 SecOrient: 1257.00 19.99 -10.01 -100.44 176.56 ---=== PROBLEMS IN DETAIL ===---